### v1

# Are nonlocal Lagrangian systems fatally unstable?

Abstract
We refute the false belief that Lagrangians, both higher-order and nonlocal, lead inexorably to instability, and an extensive variety of counterexamples are provided for both higher-order and genuinely nonlocal systems.
1 Introduction
The objective of this paper is to address and highlight a common misunderstanding related to the so-called “Ostrogradsky instability” and “Ostrogradsky theorem”. Despite appearing to be a minor misconception not warranting a dedicated article, its persistent prevalence and widespread occurrence have compelled us to invest effort in discussing it. A search for “Ostrogradsky theorem” typically yields two primary results. The first is the well-known Divergence Theorem, also known as the Gauss-Ostrogradsky Theorem, whose significance is unquestionable. The second result, which is the main focus of our extensive commentary here, relates to the Ostrogradsky Mémoire [1]. According to one interpretation [2, 3, 4], it states:
“A non-degenerate Lagrangian dependent on higher-than-first-order time derivatives corresponds to a Hamiltonian that is unbounded from below.”
In contrast, another interpretation [5, 6, 7, 8] suggests:
“In any non-degenerate theory with a fundamental dynamical variable of higher than second-order time derivative, there exists a linear instability.”
The first interpretation indeed represents a theorem, albeit a relatively straightforward one. The latter, however, is incorrect and thus cannot be considered a theorem, as we will demonstrate through various counterexamples. In this context, several authors [2, 4, 6, 7, 8, 9, 10] refer to the “Ostrogradsky instability.” Notably, there is no documentary evidence directly linking Ostrogradsky to these assertions, although his Mémoire [1] is frequently cited. A thorough examination of this source reveals no such statements. However, obtaining the document is not an easy task, and it is cumbersome to read for a contemporary scholar because of the archaic notation, more closely related to that of Lagrange and Poisson [12, 13] than to the notation of Jacobi and Hamilton [14, 15], which has ended up being imposed [16]. Ostrogradsky does not refer to the “Hamiltonian” —he does not mention the prior work of Hamilton and Jacobi, though he declares that he knows the existence of a letter from the latter, which he says he has not read. Moreover, Ostrogradsky’s work does not address the function of the Hamiltonian in this context, nor the notion of being unbounded from below.
In our view, inferring that a system is unstable from the unboundedness of the Hamiltonian is nothing but an instance of the so-called fallacy of the converse or affirming the consequent [11]. While the presence of a strict energy minimum is a sufficient condition for stability (as per the Lagrange-Dirichlet theorem [12]), it is not a necessary condition. We will elaborate on this in Section 2, where the main definitions, theorems and sufficient conditions on stability are outlined. In Section 3 these notions are applied to the second-order Pais-Uhlenbeck oscillator, either free [17] or interacting [18]. We do not consider specifically higher-order PU oscillators because the treatment is very similar. In Section 4.1 we study the stability of some genuinely nonlocal systems, particularly the infinite-order Pais-Uhlenbeck system and an instance of half-retarded/half-advanced oscillator.
In almost all examples we study, Liapunov’s corollary is used as a sufficient condition indicating stability, in spite of the fact that the Hamiltonian is not positive definite.
2 The theory of stability: a brief outline
Here we outline the main notions of stability of an equilibrium position as presented in ref. [19] from Liapunov’s work [20].
Definition 1 (Lipaunov stability)
Let $q_{j}=\dot{q}_{i}=0\,,\;\,i,j=1\ldots n$ be an equilibrium position of a dynamical system. It is said to be stable if for “small enough” initial deviations, $\,q_{j0}\,,\;\dot{q}_{i0}\,$, the system does not leave a “small” neighborhood of the equilibrium position, that is $\forall\;\varepsilon>0\,$, it exists $\delta(\varepsilon)>0\,$ such that
$\overline{\rm ext\,}\{|q_{j0}|,|\dot{q}_{i0}|\}_{i,j=1...n}<\delta\qquad\Rightarrow\qquad\overline{{\rm ext\,}}\{|q_{j}(t)|,|\dot{q}_{i}(t)|\}_{i,j=1...n}<\varepsilon\,,\quad\forall\;t>0\,.$
The equilibrium position is said to be asymptotically stable if the “small enough” initial deviation tends to vanish for $\,t\rightarrow\infty\,$, that is
$\overline{{\rm ext\,}}\{|q_{j0}|,|\dot{q}_{i0}|\}_{i,j=1...n}<\delta\qquad\Rightarrow\qquad\lim_{t\rightarrow\infty}q_{j}(t)=0\,.$
There is a variety of theorems concerning the stability of an equilibrium position. The most popular among them is [12]:
Theorem 1 (Lagrange)
If the potential energy of a conservative system has a strict minimum at a given position, then this is a stable equilibrium point [^1] It was rigorously proved by Dirichlet [21].
Since kinetic energy is generally non-negative, an immediate Corollary follows from replacing the potential energy with the energy, $E\,$.
Later Liapunov [22] generalized Lagrange theorem by replacing $\,E\,$ with a function $\,V(q,\dot{q})\,$ that: (a) is continuously derivable, (b) has a strict minimum at the equilibrium position and (c) does not increase under any motion whatsoever, namely
Theorem 2 (Liapunov)
Let $q_{j}=\dot{q}_{i}=0\,,\quad i,j=1\ldots n$ be an equilibrium position of an autonomous system and let $\,V(q,\dot{q})\,$ be a class $\mathcal{C}^{1}$ function which has a strict minimum at $\,q_{j0}=\dot{q}_{i0}=0\,$, whereas its total time derivative, $\,\dot{V}(q,\dot{q},\ddot{q})\,$, has a maximum. Then the equilibrium is stable. Furthermore, if the maximum of $\dot{V}$ is an strict maximum, then the equilibrium position is asymptotically stable.
The function $\,V(q,\dot{q})\,$ is usually referred to as Liapunov function. A particular instance of Theorem 2 is the following:
Corollary 1 (Liapunov)
If $\,V(q,\dot{q})\,$ is an integral of motion and it has a strict minimum at $q_{j}=\dot{q}_{i}=0\,,\,\,i,j=1\ldots n$, then it is a stable equilibrium point.
Both theorems, and the corollary too, establish sufficient, but not necessary, conditions for stability. Setting general necessary conditions is more difficult, except for linear systems [23] and even more so if they have constant coefficients, as
$\dot{\mathbf{X}}=\mathbb{M}\,\mathbf{X}\,,\qquad{\rm where}\qquad\mathbf{X}=\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\end{array}\right)\,,$ (1)
$\,\mathbb{M}\,$ is a constant square $n\times n$ matrix and the variables $x_{j}$ stand for the coordinates $q$ and velocities $\dot{q}\,$ of the system. It has one equilibrium point, namely $\,x_{k}=0\,,\quad k=1\ldots n\,$
Theorem 3
(a) If the real parts of all characteristic roots of $\,\mathbb{M}$ are negative, then the equilibrium point is asymptotically stable and (b) if $\mathbb{M}$ has a characteristic root whose real part is positive, then it is unstable [23].
Next theorem considers the case that some among the characteristic roots are imaginary.
Theorem 4
(a) If the characteristic roots of $\mathbb{M}$ either have negative real part or are imaginary and simple, then the equilibrium point is stable and (b) if either the real part of one among the characteristic roots is positive or one of the characteristic roots is imaginary and multiple, then it is unstable.
See Appendix A for the details of the proof.
3 The Pais-Uhlenbeck oscillator
We study the second-order Pais-Uhlenbeck oscillator in three different situations. First the free case, then interacting with an external oscillator and finally the case of an internal interaction between the two degrees of freedom of the oscillator is considered. Our aim is to show that a large class of values of the coupling constant lead to a stable behaviour, despite the fact that the Lagrangian contains second-order derivatives.
3.1 The free Pais-Uhlenbeck oscillator.
The free Pais-Uhlenbeck oscillator [17] is ruled by the fourth order Lagrangian
$L=-\frac{1}{2}\,q\,\left(1+\frac{D^{2}}{\omega_{1}^{2}}\right)\,\left(1+\frac{D^{2}}{\omega_{2}^{2}}\right)\,q\,,\qquad\mathrm{with}\qquad D:=\frac{{\rm d}}{{\rm d}t}\,,$ (2)
which is equivalent to the second-order Lagrangian
$L_{0}=-\frac{1}{2}\,q^{2}+\frac{1}{2}\,\left(\frac{1}{\omega_{1}^{2}}+\frac{1}{\omega_{2}^{2}}\right)\dot{q}^{2}-\frac{1}{2\omega_{1}^{2}\omega_{2}^{2}}\,\ddot{q}^{2}\,,\qquad\omega_{1}^{2}<\omega_{2}^{2}\,,$ (3)
apart from boundary terms. To set up the Hamiltonian formalism we implement the Ostrogradskii procedure and write $\;X_{1}:=q\,,\quad X_{2}:=\dot{q}\,,$ whereas the respective conjugate momenta are
$P_{2}$ ::=:= $\frac{\partial L}{\partial\ddot{q}}=-\frac{1}{\omega_{1}^{2}\omega_{2}^{2}}\ddot{q}\,,\hskip 95.00014pt\ddot{q}=-\omega_{1}^{2}\omega_{2}^{2}\,P_{2}$ (4)
$P_{1}$ ::=:= $\frac{\partial L}{\partial\dot{q}}-\frac{{\rm d}P_{2}}{{\rm d}t}=\frac{(\omega_{1}^{2}+\omega_{2}^{2})\,\dot{q}+\stackrel{{\scriptstyle\ldots}}{{q}}}{\omega_{1}^{2}\omega_{2}^{2}}\,,\qquad\quad\;\stackrel{{\scriptstyle\ldots}}{{q}}=\omega_{1}^{2}\omega_{2}^{2}\,P_{1}-(\omega_{1}^{2}+\omega_{2}^{2})\,X_{2}\,.$ (5)
The non-vanishing elementary Poisson brackets are $\;\{X_{a},P_{a}\}=1\,,\;\,a=1,2\,,$ and the Hamiltonian is
$H=P_{1}X_{2}-\frac{\omega_{1}^{2}\omega_{2}^{2}}{2}\,P_{2}^{2}-\frac{\omega_{1}^{2}+\omega_{2}^{2}}{2\omega_{1}^{2}\omega_{2}^{2}}\,X_{2}^{2}+\frac{1}{2}\,X_{1}^{2}\,.$ (6)
Now, following ref. [17] we introduce the new coordinates and momenta
$q_{1}=\frac{\omega_{2}\,\left(X_{1}-\omega_{1}^{2}P_{2}\right)}{\omega_{1}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,\hskip 5.0pt$ $q_{2}=\frac{\omega_{1}\,\left(X_{1}-\omega_{2}^{2}P_{2}\right)}{\omega_{2}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,$ (7)
$p_{1}=-\frac{\omega_{1}\,\left(X_{2}-\omega_{2}^{2}P_{1}\right)}{\omega_{2}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,$ $p_{2}=\frac{\omega_{2}\,\left(X_{2}-\omega_{1}^{2}P_{1}\right)}{\omega_{1}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}$ (8)
that are also a canonical set,
$\{q_{a},q_{b}\}=0\,,\qquad\{q_{a},p_{b}\}=\delta_{ab}\,,\qquad\{p_{a},p_{b}\}=0\,,$
in terms of which the Hamiltonian becomes
$H=\frac{1}{2}\,\left(p_{1}^{2}+\omega_{1}^{2}q_{1}^{2}\right)-\frac{1}{2}\,\left(p_{2}^{2}+\omega_{2}^{2}q_{2}^{2}\right)\,.$ (9)
Stability
Since the Hamiltonian is the result of two terms with opposite signs, it has no definite sign and we cannot infer the stability of the equilibrium position $\,q_{a}=p_{b}=0\,$ from the Lagrange theorem. However, other theorems, from among those listed in Section 2, will show that it is stable.
The Hamilton equations are
$\dot{q}_{1}=p_{1}\,,\qquad\dot{q}_{2}=-p_{2}\,,\qquad\dot{p}_{1}=-\omega_{1}^{2}\,q_{1}\,,\qquad\dot{p}_{2}=\omega_{2}^{2}\,q_{2}\,,$
whence it follows that
$\ddot{q}_{a}+\omega_{a}^{2}\,q_{a}=0\,,\qquad a=1,2\,,$ (10)
that is two harmonic oscillators with frequencies $\omega_{a}\,$, which are stable.
Although the Hamiltonian has no definite sign, there are other integrals of motion, particularly the energy of each oscillator:
$h_{a}:=\frac{1}{2}\,\left(p_{a}^{2}+\omega_{a}^{2}q_{a}^{2}\right)\,,\qquad a=1,2\,.$
The separate conservation of these prevents the energy transfer from one degree of freedom to the other. To put it in terms of Corollary 1, $\,G:=h_{1}+h_{2}\,$ is a positive definite integral of motion that has a strict minimum at the equilibrium point $\,q_{a}=p_{b}=0\,$, and therefore it is stable, though not asymptotically so.
3.2 The Pais-Uhlenbeck oscillator with an external interaction.
Now consider a Pais-Uhlenbeck oscillator (3) interacting with an external harmonic oscillator $\,L_{1}=\frac{1}{2}\,m\,\left(\dot{x}^{2}-\omega_{0}^{2}x^{2}\right)\,$ through an interaction Lagrangian $\;L_{\rm int}=g\,\dot{q}\,x\,$. This is a toy model for a “monochromatic electromagnetic field” $x$ interacting with a charged PU oscillator. The total Lagrangian is
$L=-\frac{1}{2}\,q^{2}+\frac{1}{2}\,\left(\frac{1}{\omega_{1}^{2}}+\frac{1}{\omega_{1}^{2}}\right)\dot{q}^{2}-\frac{1}{2\omega_{1}^{2}\omega_{1}^{2}}\,\ddot{q}^{2}+\frac{1}{2}\,m\,\left(\dot{x}^{2}-\omega^{2}x^{2}\right)+g\,\dot{q}\,x\,,$ (11)
which is second order with respect to $q$ and first order with respect to $x$. To implement the Ostrogradski procedure, we write $\;X_{1}:=q\,,\;\,X_{2}:=\dot{q}\;\,{\rm and}\;\,X_{0}:=x\,$, and the respective conjugate momenta are
[…]
$P_{1}$ ::=:= $\frac{\omega_{1}^{2}+\omega_{2}^{2}}{\omega_{1}^{2}\omega_{2}^{2}}\,\dot{q}+\frac{1}{\omega_{1}^{2}\omega_{2}^{2}}\,\stackrel{{\scriptstyle\ldots}}{{q}}+gx\,,\quad\;\stackrel{{\scriptstyle\ldots}}{{q}}=\omega_{1}^{2}\omega_{2}^{2}\,P_{1}-(\omega_{1}^{2}+\omega_{2}^{2})\,X_{2}-gX_{0}$ (13)
$P_{0}$ ::=:= $m\,\dot{x}\,,\hskip 110.00017pt\dot{x}=\frac{P_{0}}{m}\,.$ (14)
The non-vanishing Poisson brackets are $\;\{X_{a},P_{a}\}=1\,,\;\,a=0,1,2\,,$ and the Hamiltonian is
$H=P_{1}X_{2}-\frac{\omega_{1}^{2}\omega_{2}^{2}}{2}\,P_{2}^{2}+\frac{P_{0}^{2}}{2m}-\frac{\omega_{1}^{2}+\omega_{2}^{2}}{2\omega_{1}^{2}\omega_{2}^{2}}\,X_{2}^{2}+\frac{1}{2}\,X_{1}^{2}+\frac{m\omega_{0}^{2}}{2}\,X_{0}^{2}-g\,X_{2}\,X_{0}\,.$ (15)
Similar to Section 3.1, we introduce the Pais-Uhlenbeck coordinates and momenta
$q_{0}=X_{0}\,,\qquad q_{1}=\frac{\omega_{2}\,\left(X_{1}-\omega_{1}^{2}P_{2}\right)}{\omega_{1}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,\qquad q_{2}=\frac{\omega_{1}\,\left(X_{1}-\omega_{2}^{2}P_{2}\right)}{\omega_{2}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,$ (16)
$p_{0}=P_{0}\,,\qquad p_{1}=-\frac{\omega_{1}\,\left(X_{2}-\omega_{2}^{2}P_{1}\right)}{\omega_{2}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,\qquad p_{2}=\frac{\omega_{2}\,\left(X_{2}-\omega_{1}^{2}P_{1}\right)}{\omega_{1}\,\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,,$ (17)
whose Poisson brackets are ($\,a,b=0,1,2\,$):
$\{q_{a},q_{b}\}=0\,,\qquad\{q_{a},p_{b}\}=\delta_{ab}\,,\qquad\{p_{a},p_{b}\}=0\,.$
In terms of these coordinates the Hamiltonian reads
$H=\frac{1}{2}\,\left(p_{1}^{2}+\omega_{1}^{2}q_{1}^{2}\right)-\frac{1}{2}\,\left(p_{2}^{2}+\omega_{2}^{2}q_{2}^{2}\right)+\frac{p_{0}^{2}}{2m}+\frac{m\omega_{0}^{2}}{2}\,q_{0}^{2}+H_{\rm int}\,,$ (18)
where
$H_{\rm int}=-\hat{g}q_{0}\,(p_{1}+p_{2})\,,\qquad{\rm with}\qquad\hat{g}:=\frac{g\,\omega_{1}\omega_{2}}{\sqrt{\omega_{2}^{2}-\omega_{1}^{2}}}\,.$
Although the Hamiltonian has not a definite sign, we will show that the system is stable for a wide range of values of the coupling constant $g$. The Hamilton equations are:
$\dot{q}_{1}=p_{1}-\hat{g}q_{0}\qquad$ $\qquad\dot{q}_{2}=-p_{2}-\hat{g}q_{0}\qquad$ $\qquad\dot{q}_{0}=\frac{p_{0}}{m}$ (19)
$\dot{p}_{1}=-\omega_{1}^{2}\,q_{1}\qquad$ $\qquad\dot{p}_{2}=\omega_{2}^{2}\,q_{2}\qquad$ $\qquad\dot{p}_{0}=-m\omega_{0}^{2}q_{0}+\hat{g}\,(p_{1}+p_{2})\,.$ (20)
This is a linear differential system that can be written in matrix form as
$\frac{{\rm d}\mathbf{X}}{{\rm d}t}=\mathbb{M}\,\mathbf{X}\,,$ (21)
where the matrix and the vector respectively are
$\mathbb{M}=\left(\begin{array}[]{ccc|ccc}0&0&-\hat{g}&1&0&0\\ 0&0&-\hat{g}&0&-1&0\\ 0&0&0&0&0&1/m\\ \hline\cr-\omega_{1}^{2}&0&0&0&0&0\\ 0&\omega_{2}^{2}&0&0&0&0\\ 0&0&-m\omega_{0}^{2}&\hat{g}&\hat{g}&0\end{array}\right)\qquad{\rm and}\qquad\mathbf{X}=\left(\begin{array}[]{c}q_{1}\\ q_{2}\\ q_{0}\\ p_{1}\\ p_{2}\\ p_{0}\end{array}\right)\,.$ (22)
The characteristic determinant is
$\chi_{\mathbb{M}}(\lambda):=\det(\mathbb{M}-\lambda\mathbb{I})\equiv(\omega_{0}^{2}+\lambda^{2})(\omega_{1}^{2}+\lambda^{2})(\omega_{2}^{2}+\lambda^{2})+\gamma\,\lambda^{2}\,,$
with
$\,\gamma:=\frac{\hat{g}^{2}(\omega_{2}^{2}-\omega_{1}^{2})}{m}=\frac{g^{2}\,\omega_{1}^{2}\omega_{2}^{2}}{m}>0\,.$
As seen in Theorems 3 and 4 in Section 2, the stability of the system is indicated by the kind of characteristic roots. Since $\chi_{\mathbb{M}}(\lambda)$ is an even function of $\lambda$, the roots come in pairs, ±$\pm\lambda_{j}$, and the condition $\,{\mathbb{R}}{\rm e}(\pm\lambda_{j})\leq 0\,,\;\,\forall\lambda_{j}\,$, amounts to $\,{\mathbb{R}}{\rm e}(\pm\lambda_{j})=0\,$ or, equivalently, $\,\lambda_{j}^{2}<0\,$. (The case $\,\lambda_{j}=0\,$ is excluded because it would imply that 0 is a double root of $\chi_{\mathbb{M}}(\lambda)\,$, which carries instability.)
This means that, if the polynomial $\,P(\xi):=(\omega_{0}^{2}+\xi)(\omega_{1}^{2}+\xi)(\omega_{2}^{2}+\xi)+\gamma\,\xi\;$ has three simple negative roots, then the system (21) is stable.
Since $\gamma>0\,$, $\,P(\xi)>0\,$ for all real non-negative $\,\xi\,$ and therefore any real root must be negative. Thus it remains to be seen that all roots of the polynomial are real and simple. We will prove it with the help of the Implicit Function Theorem [24]. Consider the function
$F(\xi,\gamma)=(\omega_{0}^{2}+\xi)(\omega_{1}^{2}+\xi)(\omega_{2}^{2}+\xi)+\gamma\,\xi\,,$
[…]
By the Implicit Function Theorem, for each $j$, two neighbourhoods, $\,\mathcal{V}_{j}\ni 0\,$ and $\,\mathcal{U}_{j}\ni-\omega_{j}^{2}\,$, and a differentiable function $\;f_{j}:\mathcal{V}_{j}\rightarrow\mathcal{U}_{j}\,$ exist such that
$f_{j}(0)=-\omega_{j}^{2}\qquad{\rm and}\qquad F(f_{j}(\gamma),\gamma)=0\,,\qquad\gamma\in\mathcal{V}_{j}\,.$
Since the functions $f_{j}$ are continuous, an open subset $\,\mathcal{V}\,$ exists such that
$\mathcal{V}\subset\mathcal{V}_{j}\,,\quad\forall j=0,1,2\,,\qquad{\rm and}\qquad f_{j}(\mathcal{V})\cap f_{k}(\mathcal{V})=\emptyset\,,\quad j\neq k\,.$
Hence, for all $\,\gamma\in\mathcal{V}\,,\quad f_{j}(\gamma)\neq f_{k}(\gamma)\,,\quad j\neq k\,$, and therefore $\,f_{j}(\gamma)\,,\;\,j=0,1,2\,$ are three real simple roots of the polynomial, indeed
$P(f_{j}(\gamma))\equiv F(f_{j}(\gamma),\gamma)=0\,.$
3.3 Pais-Uhlenbeck oscillator with an internal interaction
Are there internal interactions connecting both oscillators of a PU system such that stability is preserved? Let
$H=h_{1}-h_{2}+V(q,p)\,,\qquad h_{a}=\frac{1}{2}\,\left(p_{a}^{2}+\omega_{a}^{2}q_{a}^{2}\right)\,$ (23)
be a perturbation of the PU Hamiltonian (9). We look for a Liapunov function $G(q,p)\,$, that is (a) an integral of motion $\;\{H,G\}=0\,$ and (b) has a strict minimum at $q_{a}=p_{b}=0\,$.
Writing the coordinates as $\,(x_{i})=(q_{a},p_{b})\,$, with $i=1\ldots 4\,$, the condition that $G$ is an integral of motion implies that
$\partial_{i}H\,\Omega_{ij}\,\partial_{j}G=0\,\qquad{\rm or}\qquad\nabla H^{T}\,\Omega\,\nabla G=0\,,$ (24)
where summation over repeated indices is understood and
$\Omega=\left(\begin{array}[]{cc|cc}0&0&1&0\\ 0&0&0&1\\ \hline\cr-1&0&0&0\\ 0&-1&0&0\end{array}\right)\,.$
In the case that $V$ is a homogeneous quadratic function, namely ${V=\frac{1}{2}\,V_{ij}x_{i}x_{j}}\,$ [^2] Summation over repeated indices is understood unless otherwise stated, the Hamiltonian is quadratic too and equation (24) admits a particular solution that is also quadratic ${G=\frac{1}{2}\,G_{ij}x_{i}x_{j}}\,$. Obviously, $\,V_{ij}=V_{ji}\,$, $\,G_{ij}=G_{ji}\,$ and $\,H_{ij}=H_{ji}\,$, and the equation reads
$H_{il}\,\Omega_{ij}\,G_{jk}\,x_{l}x_{k}=0\,,\qquad\forall x_{l}\,,$
which amounts to
$H_{li}\,\Omega_{ij}\,G_{jk}+H_{ki}\,\Omega_{ij}\,G_{jl}=0\qquad{\rm or}\qquad\mathbb{H}\,\Omega\,\mathbb{G}+\left(\mathbb{H}\,\Omega\,\mathbb{G}\right)^{T}=0\,,$ (25)
where the square 4$\times$4 matrices $\;\mathbb{H}=\left(H_{ij}\right)\,$ and $\;\mathbb{G}=\left(G_{ij}\right)\,$ have been introduced. In the present quadratic case, the strict minimum condition is equivalent to require the positivity of $\,G(x)\,$ for $x_{j}\neq 0\,$.
The case $\,V=V(q_{a})\,$ is solved in Appendix B, where we show that a Liapunov function exists as long as the coefficients of the interaction potential fulfill the inequalities (72),
$\left(a+b+\omega_{1}^{2}-\omega_{2}^{2}\right)^{2}>4c^{2}>4\,\left(a+\omega_{1}^{2}\right)\,\left(b-\omega_{2}^{2}\right)\qquad{\rm and}\qquad\left|b-\omega_{2}^{2}\right|>\left|a+\omega_{1}^{2}\right|\,$ (26)
and the Liapunov function is (73)
$G=h_{1}+h_{2}+V+\frac{1}{v+w}\,\left(c^{2}q_{1}^{2}-\left[c^{2}+b(a+b+\omega_{1}^{2}-\omega_{2}^{2})\right]\,q_{2}^{2}+2c\left[b-\omega_{2}^{2}\right]q_{1}q_{2}+2cp_{1}p_{2}\right)\,.$
Therefore Liapunov’s Corollary 1 ensures the stability of the system with the interaction potential $V(q)$ whenever the inequalities (26) are satisfied.
3.4 A double oscillator with ghosts without runaway instabilities
We shall now study the stability of the double-oscillator Hamiltonian (23), with $\omega_{1}=\omega_{2}=1$, perturbed by an internal interaction term. The system is not a perturbed Pais-Uhlenbeck oscillator because the latter requires $\omega_{1}\neq\omega_{2}$, otherwise the transformations (7-8) do not apply. However, its interest lies in extending the proposal of [18], which we shall work in a systematic fashion that relies on Liapunov’s Corollary 1.
The perturbed Hamiltonian is
$H=h_{1}-h_{2}+V(q)\,,\qquad h_{a}=\frac{1}{2}\,\left(p_{a}^{2}+q_{a}^{2}\right)\,,$ (27)
and we look for an integral of motion
$G=h_{1}+h_{2}+k^{2}+W(q)\,,\qquad{\rm with}\qquad k=p_{1}q_{2}+p_{2}q_{1}\,.$ (28)
The condition $\,\{H,G\}=0\,$ easily leads to
$p_{1}\,\left(-W_{1}+V_{1}+2\,q_{2}\,\left[q_{2}V_{1}+q_{1}V_{2}\right]\right)+p_{2}\,\left(W_{2}+V_{2}+2\,q_{1}\,\left[q_{2}V_{1}+q_{1}V_{2}\right]\right)=0\,,$
where ${W_{b}:=\frac{\partial W}{\partial q^{b}}\,}$ and so on. Since it must hold for all values of $p_{b}$, it implies that
$W_{1}=\left(1+2q_{2}^{2}\right)\,V_{1}+2q_{1}q_{2}V_{2}\qquad{\rm and}\qquad-W_{2}=\left(1+2q_{1}^{2}\right)\,V_{2}+2q_{1}q_{2}V_{1}\,.$ (29)
The solutions of this equation are analysed in Appendix C for the particular case $\,V=V(z)\,$, where $\,z(q_{a})\,$ is a quartic polynomial. There are four classes of $\,z(q_{a})\,$ for which the equation (29) admits a solution and are given by equations (77-80), namely
  • I.-
    $\,\displaystyle{V=\frac{C}{1-b^{2}+2\,q_{1}^{2}(1+b)+2\,q_{2}^{2}(1-b)}\qquad{\rm and}\qquad W=\frac{1}{b}\,\left(1+2q_{1}^{2}+2q_{2}^{2}\right)\,V}\,,$
  • II.-
    $\,\displaystyle{V=C\,\left[2\,\left(q_{1}^{2}-q_{2}^{2}\right)+a\right]\qquad{\rm and}\qquad W=C\,\left(1+2q_{1}^{2}+2q_{2}^{2}\right)}\,,$
  • III.-
    $\,\displaystyle{V=\frac{C}{\sqrt{1+2\,(q_{1}^{2}+q_{2}^{2})+(q_{1}^{2}-q_{2}^{2})^{2}}}\qquad{\rm and}\qquad W=\left(q_{2}^{2}-q_{1}^{2}\right)V}\,,$
  • IV.-
    $\,\displaystyle{V=\frac{C}{1+2\,(q_{1}^{2}+q_{2}^{2})}\qquad{\rm and}\qquad W=2\left(q_{2}^{2}-q_{1}^{2}\right)V}\,.$
In each case, according to Corollary 1, the local strict minima of the integral of motion
$G=\frac{1}{2}\,\left(p_{1}^{2}+q_{1}^{2}+p_{2}^{2}+q_{2}^{2}\right)+\left(p_{1}q_{2}+p_{2}q_{1}\right)^{2}+W(q)$
are points of stable equilibrium. After a little algebra, which is simple but tedious, we obtain that for all equilibrium points $p_{1}=p_{2}=0\,$ and, as for coordinates, the different cases are listed in the following table:
I(A) $q_{1}=0$ $q_{2}=0$ ${(b-1)^{2}>\frac{2C}{(b+1)}\quad{\rm and}\quad(b+1)^{2}>\frac{2C}{(b-1)}}$
I(C) $q_{1}=0$ ${q^{2}_{2}=\sqrt{\frac{C}{b-1}}-\frac{b+1}{2}}$ ${\sqrt{\frac{C}{b-1}}\geq\frac{b+1}{2}\,,\quad b<1\,,\quad C<0}$
I(D) $q_{2}=0$ ${q^{2}_{1}=\sqrt{\frac{2C}{b+1}}-\frac{1-b}{2}}$ ${\sqrt{\frac{C}{b+1}}\geq\frac{1-b}{2}\,,\quad b>-1\,,\quad C>0}$
II(A) $q_{1}=0$ $q_{2}=0$ $C>-\frac{1}{4}\,$
III(A) $q_{1}=0$ $q_{2}=0$ $|C|<\displaystyle{\frac{1}{2}}\,$
III(C) $q_{1}=0$ $q^{2}_{2}=\sqrt{-2C}-1\,$ $C\leq-\displaystyle{\frac{1}{2}}\,$
III(D) $q_{2}=0$ $q_{1}=\sqrt{2C}-1$ $C\geq\displaystyle{\frac{1}{2}}\,$
IV(A) $q_{1}=0$ $q_{2}=0$ $|C|<\displaystyle{\frac{1}{4}}\,$
IV(C) $q_{1}=0$ $q^{2}_{2}=\sqrt{-2C}-1\,$ $C\leq-\displaystyle{\frac{1}{4}}\,$
IV(D) $q_{2}=0$ $q^{2}_{1}=\sqrt{2C}-1$ $C\geq\displaystyle{\frac{1}{4}}\,$
Recall that Liapunov’s Corollary is only a sufficient condition.
Actually case III is the one studied in ref. [18], where it is proved that, in spite of the presence of ghosts, after a small perturbation of the equilibrium, the system remains confined in a region around the origin whose size is a multiple of $|C|\,$. However confinement does not mean stability: the system could be lead away to the “wall” of the confinement area, no matter how small is the initial perturbation. They also discuss stability by linearizing the equations near the equilibrium point. The method we present here is more powerful in that it permits to exactly conclude the stability of the equilibrium point $q_{1}=q_{2}=0\,$ as far as the coupling constant meets $|C|<1/2\,$.
4 Genuine nonlocallity.
So far we have considered a finite-order Pais-Uhlenbeck oscillator —actually only second-order Lagrangians—but for higher-orders the technique is the same and the results are very similar. We will now study systems that are genuinely nonlocal, namely (i) a PU oscillator of infinite order and (ii) a harmonic oscillator self-interacting with its half-advanced/half-retarded configurations.
Whereas in the local case Lagrange equations form an ordinary differential system of $2n^{\rm th}$-order, in the nonlocal case the equations contain some terms that depend on whole fragments of the trajectory. As a consequence, while in the local case a finite number of initial data, i. e. the coordinates and their derivatives up to order $12n-$ 1, are enough to determine the future evolution, for nonlocal equations it is not so. As a rule, an infinite number of data $\,(\alpha_{j})_{j\in\mathcal{J}}\,$ are needed to determine a solution, and often these data do not belong to a fixed instant of time. With this in mind, Definition 1 for stability must be adapted to nonlocal equations of motion.
Definition 2
Given a nonlocal system having an equilibrium position $\,q_{0}\,$ that corresponds to the data $\,(\alpha_{0j})_{j\in\mathcal{J}}\,$, that is $q(t,\alpha_{0j})=q_{0}\,$, this equilibrium position is said stable if, $\forall\varepsilon>0\,$, it exists $\delta>0$ such that $\|\alpha_{j}-\alpha_{0j}\|<\delta\,$ implies that $\left|q(t,\alpha_{j})-q_{0})\right|<\delta\,$, $\forall t\,$, for some given norm in the space of data.
4.1 Fully nonlocal Pais-Uhlenbeck oscillator
When the operator $F(D)$ in the Lagrangian (2) is an infinite product,
$L=-q\,F(D)q\,,\qquad\quad F(D)=\prod_{n=1}^{\infty}\left(1+\frac{D^{2}}{\omega_{n}^{2}}\right)\,,$ (30)
then it is completely nonlocal. For the reasons given in [27], instead of dealing with an infinite number of derivatives, here we will adopt the functional formalism for nonlocal systems and convert the operator $F(D)$ into an integral operator through the Fourier transform
$F(D)q(t)$ ::=:= $\frac{1}{\sqrt{2\pi}}\,F(D)\int_{\mathbb{R}}{\rm d}k\,e^{-ikt}\,\tilde{q}(k)=\frac{1}{\sqrt{2\pi}}\,\int_{\mathbb{R}}{\rm d}k\,F(D)e^{-ikt}\,\tilde{q}(k)$
$=$ $\frac{1}{\sqrt{2\pi}}\,\int_{\mathbb{R}}{\rm d}k\,\prod_{n=1}^{\infty}\left(1+\frac{(-ik)^{2}}{\omega_{n}^{2}}\right)\,\,e^{-ikt}\,\tilde{q}(k)=\int_{\mathbb{R}}{\rm d}\tau\,K(t-\tau)\,q(\tau)\,.$
We thus have that
$F(D)q=K\ast q\,,\qquad{\rm with}\qquad K(t):=\frac{1}{2\pi}\,\int_{\mathbb{R}}{\rm d}k\,F(-ik)\,e^{-ikt}\,.$ (31)
The function in the integral is defined by the infinite product
$F(-ik):=\prod_{n=1}^{\infty}\left(1-\frac{k^{2}}{\omega_{n}^{2}}\right)\,.$
A sufficient condition for the product to be absolutely convergent is that the series $\,\displaystyle{\sum_{n=1}^{\infty}\frac{1}{\omega_{n}^{2}}}\,$ is convergent —see [25]. In this case the infinite product converges uniformly on any disk $\,|k|<R\,$ in the complex plane and $F(z)$ is an entire function of the complex variable $z\,.$
We will consider in detail the particular case $\,\omega_{n}=\pi\,n\,$ and use the infinite product [25]
$F(z)=\prod_{n=1}^{\infty}\left(1+\frac{z^{2}}{\omega_{n}^{2}}\right)=\frac{\sinh z}{z}=\sum_{n=0}^{\infty}\frac{z^{2n}}{(2n+1)!}\,.$
Applying the definition (31), we have that
$K(t):=\frac{1}{2\pi}\,\int_{\mathbb{R}}{\rm d}k\,\frac{\sin k}{k}\,e^{-ikt}=\frac{1}{2}\,\Theta(1-|t|)\,,$ (32)
hence $F(D)$ is the convolution operator
$F(D)q=K\ast q\,,\qquad{\rm with}\qquad K(t)=\frac{1}{2}\,\Theta(1-|t|)\,.$ (33)
The Lagrange equation is $\,K\ast q=0\,$ and, including (32), it can be written as
$\int_{-1}^{1}{\rm d}\tau\,q(t+\tau)=0\,,\qquad\forall t\in\mathbb{R}\,.$ (34)
$q(t)=0\,$ is of course a solution and 00 is an equilibrium point and, in order to assess its stability on the basis of Definition 2, we need to identify a convenient space of data and choose a norm on it. We may expect that $q(t)$ is a real function of class $\mathcal{C}^{0}(\mathbb{R})$, at least; hence the integral is differentiable and the first derivative of the equation is $\,q(t+1)=q(t-1)\,$. Therefore the solutions of (34) are continuous periodic functions, with period 2 and vanishing mean value, therefore each solution has a Fourier series
$q(t)\sim\sum_{n\neq 0}a_{n}\,e^{i\pi nt}\,,\qquad{\rm with}\qquad a_{n}=\frac{1}{2}\,\int_{\mathbb{R}}{\rm d}t\,q(t)\,e^{-i\pi nt}\,.$ (35)
Since $q(t)$ is real, $\,a_{-n}=a_{n}^{\ast}\,$, where means “complex conjugate”. We will refer to this class of solutions as $\,\mathcal{D}_{0}\,$.
The symbol $\sim$ is to emphasize that the sum of the series may be understood in different ways [24, 26] depending on the kind of convergence we choose, namely mean-square convergence or pointwise convergence. Since $q(t)$ is continuous in $\mathbb{R}$, it belongs to the class $L^{2}([0,2])$ and mean-square convergence is assured. However, in mathematical (classical) physics pointwise convergence is more convenient.
Since Parseval identity implies that $\,\displaystyle{\sum_{n=1}^{\infty}|a_{n}|^{2}<\infty}\,$, the second expression in (35) maps the space $\,\mathcal{D}_{0}\,$ onto the space of square summable complex sequences $\,\{a_{n}\}_{n\in\mathbb{N}}\,$.
Since for a general continuous periodic function, without further conditions, the pointwise convergence of its Fourier series is not assured, the symbol $\sim$ in the first expression (35) does not mean equality and the above map is not invertible. To pointwise recover the values of a continuous periodic function we must resort to the Fejér sums (recall $\,a_{0}=0\,$)
$S_{n}(t)=\frac{1}{n}\,\sum_{k=1}^{n-1}\,\sum_{|l|=1}^{k}a_{l}\,e^{i\pi lt}=\sum_{|l|=1}^{n-1}\left(1-\frac{|l|}{n}\right)\,a_{l}\,e^{i\pi lt}\,.$
Fejér Theorem [26] then establishes that, provided that $q(t)$ is continuous and periodic, the succession of Fejér sums $S_{n}(t)$ converges uniformly to $q(t)\,$:
$q(t)=\lim_{n\rightarrow\infty}\sum_{|l|=1}^{n-1}\left(1-\frac{|l|}{n}\right)\,a_{l}\,e^{i\pi lt}\,.$ (36)
This gives a way to recover pointwise the function from its Fourier series.
Because when we later define the Hamiltonian we will need the Fourier series of $\,\dot{q}(t)\,$, we choose the dynamic space $\mathcal{D}_{1}$ as the space of periodic $\mathcal{C}^{1}\,$ functions, with period 2 and vanishing mean. It will be coordinated by the complex sequences $\left\{a_{n}\right\}_{n\in\mathbb{N}}\,$ such that $\sum_{n=1}^{\infty}n^{2}\,|a_{n}|^{2}<\infty\,$.
We will work in the functional formalism as developed in [28] because in our view is the most convenient and we will set up a canonical formalism on the basis of a symplectic form and a Hamiltonian on the dynamic space $\mathcal{D}_{1}$, coordinated according to (35-36). Given the Lagrangian (30), the function —see [28], eq. (3)—
$\lambda(q,\tau,\sigma):=\frac{\delta L(T_{\tau}q)}{\delta q(\sigma)}=-\delta(\tau-\sigma)\,K\ast q(\tau)-q(\tau)\,K(\sigma-\tau)\,$ (37)
will be useful. The symplectic form on $\mathcal{D}_{1}$ is given by —see [28], eq. (44)—
$\omega$ $=$ $\frac{1}{2}\,\int_{\mathbb{R}^{3}}{\rm d}\tau\,{\rm d}\rho\,{\rm d}\sigma\,\frac{\delta\lambda(q,\tau,\sigma)}{\delta q(\rho)}\,\left[\Theta(\sigma)-\Theta(\rho)\right]\,\delta q(\rho)\wedge\delta q(\sigma)$
$=$ $-\int_{\mathbb{R}^{2}}{\rm d}\rho\,{\rm d}\sigma\,\left[\Theta(\sigma)-\Theta(\rho)\right]\,K(\sigma-\rho)\,\delta q(\rho)\wedge\delta q(\sigma)\,,$
where $\,K(\rho)=K(-\rho)\,$ has been included, and using that
$\int_{\mathbb{R}^{2}}{\rm d}\rho\,{\rm d}\sigma\,f(\sigma,\rho)\,\left[\Theta(\sigma)-\Theta(\rho)\right]=\int_{\mathbb{R}^{+2}}{\rm d}\rho\,{\rm d}\sigma\,\left[f(\sigma,-\rho)-f(-\sigma,\rho)\right]\,,$
it becomes
$\omega$ $=$ $\int_{\mathbb{R}^{+2}}{\rm d}\rho\,{\rm d}\sigma\,\Theta(1-[\rho+\sigma])\,\delta q(\rho)\wedge\delta q(-\sigma)$
$=$ $\int_{\mathbb{R}^{+2}}{\rm d}\rho\,{\rm d}\sigma\,\Theta(1-[\rho+\sigma])\,\lim_{n,m\rightarrow\infty}\sum_{|l|=1}^{n-1}\sum_{|j|=1}^{m-1}\left(1-\frac{|l|}{n}\right)\,\left(1-\frac{|j|}{m}\right)\,e^{i\pi\left(l\rho-j\sigma\right)}\,\delta a_{l}\wedge\delta a_{j}$
where (33) and (36) have been included, or
$\omega=\lim_{n,m\rightarrow\infty}\sum_{|l|=1}^{n-1}\sum_{|j|=1}^{m-1}\left(1-\frac{|l|}{n}\right)\,\left(1-\frac{|j|}{m}\right)\,A_{[lj]}\,\delta a_{l}\wedge\delta a_{j}\,,$
with
$A_{[lj]}:=i\,\int_{\mathbb{R}^{+2}}{\rm d}\rho\,{\rm d}\sigma\,\Theta(1-[\rho+\sigma])\,\sin\left(\pi\,\left[l\rho-j\sigma\right]\right)\,.$
Introducing the variables $\;s=\rho+\sigma\,$ and $\,\Delta=\rho-\sigma\,$, we obtain
$A_{[lj]}=\frac{i}{2}\,\int_{0}^{1}{\rm d}s\,\int_{-s}^{s}{\rm d}\Delta\,\sin\left(\pi\,\left[s\,\frac{l-j}{2}+\Delta\,\frac{l+j}{2}\right]\right)=-\frac{i}{\pi\,l}\,(-1)^{l}\,\delta_{-l\;j}\,,$
whence it follows that
$\omega=\lim_{\nu\rightarrow\infty}\sum_{l=1}^{\infty}\Omega_{l}(n,m)\,\delta a_{l}\wedge\delta a_{-l}\,,$
where we have put $\,\nu=\min\{n,m\}\,$ and
$\Omega_{l}(n,m)=\left\{\begin{array}[]{cl}\displaystyle{-\left(1-\frac{l}{n}\right)\,\left(1-\frac{l}{m}\right)\,\frac{2i}{\pi l}\,(-1)^{l}}\,,&\quad{\rm if}\;\;\nu>l\\[8.61108pt] 0\,,&\quad{\rm if}\;\;\nu<l\end{array}\right.\,.$
Taking the limit we finally arrive at
$\omega=\sum_{l=1}^{\infty}\frac{2i}{\pi l}\,(-1)^{l}\,\delta a_{l}^{\ast}\wedge\delta a_{l}\,,$ (38)
where we have used that $\,a_{-l}=a_{l}^{\ast}\,$.
In turn, the Hamiltonian restricted to $\mathcal{D}_{1}$ is —see [28], eq. (43)—
$h=\int_{\mathbb{R}^{2}}{\rm d}\rho\,{\rm d}\sigma\,\left[\Theta(\sigma)-\Theta(\rho)\right]\,\dot{q}(\sigma)\,\lambda(q,\rho,\sigma)-L(q)\,.$
The Lagrangian vanishes on $\mathcal{D}_{1}$, because the Lagrange equations $\,K\ast q=0\,$ hold on $\mathcal{D}_{1}$. As for the first term, in a similar way as we did above for the symplectic form, we obtain that
$h=2\sum_{n=1}^{\infty}(-1)^{n}\,\left|a_{n}\right|^{2}\,.$ (39)
Equations (38-39) can be written also in terms of the real variables
$P_{n}:=a_{n}+a^{\ast}_{n}\,,\qquad Q_{n}:=\frac{i(-1)^{n}}{n\pi}\,\left(a_{n}-^{\ast}_{n}\right)\,,\qquad a_{n}=\frac{1}{2}\,\left(P_{n}-i\,(-1)^{n}\,n\pi\,Q_{n}\right)\,,$ (40)
and we have that
$\omega=\sum_{n=1}^{\infty}\delta P_{n}\wedge\delta Q_{n}\qquad{\rm and}\qquad h=\sum_{n=1}^{\infty}\frac{1}{2}\,(-1)^{n}\,\left(P_{n}^{2}+n^{2}\pi^{2}\,Q_{n}^{2}\right)\,,$ (41)
which implies that the coordinates $Q_{n}$ and $P_{n}$ form a canonical conjugate pair, and the Hamiltonian has no definite sign.
Although energy is not positive, as far as the system is equivalent to an infinite set of free uncoupled oscillators, it is stable. Actually each term in the series that defines the Hamiltonian is individually preserved and therefore $\,G=\sum_{n=1}^{\infty}\,\left(P_{n}^{2}+n^{2}\pi^{2}\,Q_{n}^{2}\right)\,$ is an integral of motion that has a strict minimum at $P_{n}=Q_{m}=0\,$. Therefore, by Liapunov’s Corollary 1, the equilibrium point is stable.
4.2 The case study of a nonlocal oscillator
Consider the nonlocal Lagrangian $\,L=\displaystyle{\frac{a}{2}\,\dot{q}^{2}(t)-\frac{1}{2}\,q^{2}(t)+\kappa\,q(t+T)\,q(t)\,,\quad\kappa\neq 0}$, whose Lagrange equation is
$a\,\ddot{q}(t)+q(t)=\kappa\,\left[q(t+T)+q(t-T)\right]\,.$ (42)
We will focus on the stability, or instability, for the cases $a\neq 0$ —an oscillator coupled to its own half-retarded and half-advanced configurations— and $a=0$.
As we seek continuous solutions for all real $t$, when $a\neq 0$ they are necessarily smooth, as it follows from applying the equation and all its subsequent derivatives. If we write the equation as
$q(\tau\pm T)=-q(\tau\mp T)+\frac{1}{\kappa}\,\left[a\,\ddot{q}(\tau)+q(\tau)\right]$ (43)
it is clear that forward evolution and backward regression are both determined by the knowledge of $q(\tau)=\xi(\tau)$ in the interval $-T\leq\tau\leq T\,$. This datum is subject to the infinite number of constraints
$\xi^{(n)}(T)+\xi^{(n)}(-T)=\frac{1}{\kappa}\,\left[a\,\xi^{(n+2)}(0)+\xi^{(n)}(0)\right]\,,\qquad\forall n\in\mathbb{N}\,,$ (44)
which result from taking the limit for $\,t\rightarrow 0\,$ in equation (42) and its subsequent derivatives:
$\xi^{(n)}(T)+\xi^{(n)}(-T)=\frac{1}{\kappa}\,\left[a\,\xi^{(n+2)}(0)+\xi^{(n)}(0)\right]\,.$
Let us now introduce the sequence of functions,
[…]
Equation (42) then amounts to
$q_{n+2}(\tau)=-q_{n}(\tau)+\hat{\Theta}q_{n+1}(\tau)\,,\qquad{\rm where}\qquad\hat{\Theta}:=\frac{1}{\kappa}\,\left[a\,\frac{{\rm d}^{2}\;}{{\rm d}t^{2}}+1\right]\,,\qquad n\in\mathbb{Z}\,.$ (45)
Focussing now on the forward evolution —the backward regression works similarly— the initial datum implies the relations
$q_{0}(\tau)=\xi(\tau)\qquad{\rm and}\qquad q_{1}(\tau)=-\xi(\tau-T)+\hat{\Theta}\xi(\tau)\,.$ (46)
The evolution equation and the above relations suggest seeking solutions shaped as
$q_{n}(\tau)=X_{n}\,\xi(\tau)+Y_{n}\xi(\tau-T)\,$ (47)
that, including (45), amount to
$X_{n+2}=-X_{n}+\hat{\Theta}\,X_{n+1}\qquad{\rm with}\qquad X_{0}=1\,,\qquad X_{1}=\hat{\Theta}\,$
and
$Y_{n+2}=-Y_{n}+\hat{\Theta}\,Y_{n+1}\qquad{\rm with}\qquad Y_{0}=0\,,\qquad Y_{1}=-1\,.$
These can be solved by means of the generating functions
$F(z)=\sum_{n=0}^{\infty}X_{n}\,z^{n}\qquad{\rm and}\qquad G(z)=\sum_{n=0}^{\infty}Y_{n}\,z^{n}\,,$
and we easily obtain that
$F(z)=\frac{1}{z^{2}-\hat{\Theta}\,z+1}\qquad{\rm and}\qquad G(z)=-\frac{z}{z^{2}-\hat{\Theta}\,z+1}\,,$ (48)
whence it follows that
$Y_{n}=-X_{n-1}\,.$ (49)
To expand $F(z)$ and obtain an explicit expression for $X_{n}$, we write
$z^{2}-\hat{\Theta}\,z+1=\left(1-z\,\sigma e^{\alpha}\right)\left(1-z\,\sigma e^{-\alpha}\right)\,,\quad{\rm with}\quad\cosh\alpha=\frac{\sigma\,\hat{\Theta}}{2}\,,$
where $\;\sigma={\rm sign}(\kappa)\,$, if $a=0$, and $\,\sigma=1\,$, otherwise.
Then we have that
$F(z)=\left(1-z\,\sigma e^{\alpha}\right)^{-1}\left(1-z\,\sigma e^{-\alpha}\right)^{-1}=\sum_{m,n=0}^{\infty}z^{n+m}\,\sigma^{n+m}\,e^{(n-m)\alpha}=\sum_{r=0}^{\infty}z^{r}\,\sigma^{r}\,e^{r\alpha}\,\sum_{m=0}^{r}e^{-2m\alpha}\,,$
which easily leads to
$X_{r}=\sigma^{r}\,\frac{\sinh(r+1)\alpha}{\sinh\alpha}\qquad{\rm and}\qquad Y_{r}=-\sigma^{r-1}\frac{\sinh(r\alpha)}{\sinh\alpha}\,,\qquad\cosh\alpha=\frac{\sigma\,\hat{\Theta}}{2}\,,$ (50)
where (49) has been included.
4.2.1 The case $a=0$
Let us focus on the simplest case $a=0\,$, that is $\hat{\Theta}=1/\kappa\,$. We distinguish three cases:
Large $|\kappa|\,$.
[…]
$X_{n}=\sigma^{n}\,\frac{\sin(n+1)\omega}{\sin\omega}\qquad{\rm and}\qquad Y_{n}=-\sigma^{n-1}\,\frac{\sin(n\omega)}{\sin\omega}\,,$
and therefore $\,\left|X_{n}\right|\,$ and $\,\left|Y_{n}\right|\,$ are bounded by ${\,\frac{1}{\sin\omega}=\frac{\sqrt{4\kappa^{2}-1}}{2|\kappa|}\,}$. Then, including that $\,q(\tau+nT)=X_{n}\,\xi(\tau)-X_{n-1}\,\xi(\tau-T)\,$, we easily obtain that
$\|q\|_{\infty}\leq\frac{\sqrt{4\kappa^{2}-1}}{|\kappa|}\,\|\xi\|_{\infty}\,,\qquad{\rm where}\qquad\|q\|_{\infty}=\overline{{\rm ext}}\{|q(t)|\,,\;t\in\mathbb{R}\}\,,$
whence it follows that, provided that $\,|\xi(\tau)|<\delta\,,\;|\tau|<t\,$, then $\;\displaystyle{|q(t)|<\frac{\sqrt{4\kappa^{2}-1}}{|\kappa|}\,\delta}\,$, for all $t\,$, and the system is stable according to Definition 2.
Critical $|\kappa|=1/2\,$.
In this case, $\,\alpha=0\,$ and $\,\displaystyle{X_{n}=\sigma^{n}\,\sum_{m=0}^{n}1=\sigma^{n}\,(n+1)}\,$. Hence
$\,q(\tau+nT)=\sigma^{n}\,\left[(n+1)\,\xi(\tau)-n\,\sigma\,\xi(\tau-T)\right]=\sigma^{n}\,\left\{\xi(\tau)+n\,\left[\xi(\tau)-\sigma\,\xi(\tau-T)\right]\right\}\,$
that for suitable $\xi(\tau)\,$ might increase with $n\,$, which means that the system is unstable.
Small $|\kappa|\,$.
If $\,|\kappa|<1/2\,$, then $\,\displaystyle{X_{n}=\sigma^{n}\,\frac{\sinh(n+1)\omega}{\sinh\omega}}\,$ and therefore, for large $n$,
$|X_{n}|\sim\frac{e^{(n+1)\alpha}}{2\,|\sinh\alpha|}\,,\qquad{\rm with}\qquad\cosh\alpha=\frac{1}{2|\kappa|}\,,$
which implies that
$\|q_{n}\|_{\infty}\sim\frac{e^{(n+1)\alpha}}{2|\sinh\alpha|}\,\|\xi\|_{\infty}\,.$
Therefore
$\left|q(t)\right|\sim\frac{e^{\alpha}}{2|\sinh\alpha|}\,\|\xi\|_{\infty}\,e^{\alpha\,\left[\frac{t}{T}\right]}\,,$
where []\,[\;\,]\,[ ] means the “integer part”, and the instability of the equilibrium point $q=0$ is obvious.
4.2.2 The case $a\neq 0$
Let us go back to equation (42) with $a\neq 0\,$,
$a\,\ddot{q}(t)+q(t)=\kappa\,\left[q(t+T)+q(t-T)\right]\,,\qquad\kappa\neq 0\,.$ (51)
The equilibrium solution is $\,q(t)=0\,$, which results from the evolution of the initial datum $\,\eta(\tau)=0\,,\quad|\tau|\leq T\,$. Consider a perturbation $\,\xi(\tau)=y\,e^{i\omega\tau}\,,\quad|\tau|\leq T\,$, where $\,\omega=\omega_{1}+i\,\omega_{2}\,$ is any complex root of
$\cos(\omega T)=\frac{1-a\,\omega^{2}}{2\,\kappa}\,.$ (52)
If either $\,a\neq 0\,$ or $\,|\kappa|<1/2\,$, this equation has some roots with nonvanishing imaginary part and, since the equation is even in $\omega$, some of its roots will have $\omega_{2}<0\,$.
Let $\omega\,$ be one of these roots. It is obvious that $\,q(t)=y\,e^{i\omega t}\,$ is a solution of equation (51) and that $\,|q(t)|=|y|\,e^{-\omega_{2}t}\,,$ therefore if $\omega_{2}$ is negative, the initially small perturbation grows exponentially for large $t$ and the equilibrium point is unstable.
5 Conclusion
The purpose of this work is to refute the widely held false belief that higher-order and nonlocal Lagrangians lead fatally to instability. This belief is based on the fallacy of the converse applied to the Lagrange-Dirichlet theorem, that is taking this theorem, which is a sufficient condition for stability, as if it was a necessary condition.
We have presented several examples of higher-order and nonlocal Lagrangian systems that are stable, despite the fact that the Hamiltonian is not positive definite. We have focused mainly on variants of the second-order Pais-Uhlenbeck oscillator: free and with interaction, both internal and external, and found that the free case is always stable while the cases with interaction are stable if the coupling parameters are in a certain range. The case of higher-order oscillators can be treated in the same way and therefore we have omitted them.
Our treatment is based on the Liapunov theory and mainly on the corollary 1, which is an extension of the Lagrange-Dirichlet theorem in which the role of energy is played by an integral of motion conveniently chosen.
We have also studied some nonlocal systems, namely a PU oscillator of infinite order without interaction, which turned out to be stable, and a nonlocal oscillator model which, depending on the value of a parameter, is stable or not.
Acknowledgement
J. Ll. was supported by the Spanish MINCIU and ERDF (project ref. PID2021-123879OB-C22), and C. H. expresses gratitude to Ivan Kolá$\check{\text{r}}$ for the valuable discussions on this topic and to the Nennisiwok AI Lab team for their support.
Appendix A
Let us consider the linear system
$\dot{\mathbf{X}}=\mathbb{M}\,\mathbf{X}\,,\qquad{\rm with}\qquad\mathbf{X}=\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\end{array}\right)$ (53)
and $\,\mathbb{M}\,$ is a constant square $n\times n$ matrix. This system has one equilibrium point, namely $\,x_{k}=0\,,\quad k=1\ldots n\,$.
We aim to proof that, provided that the characteristic roots of $\mathbb{M}$ either have negative real part or are imaginary and simple, then the equilibrium point is stable. We start introducing the linear transformation
$\mathbf{X}=\mathbb{A}\,\mathbf{Y}\,,\qquad\mathbf{Y}=\mathbb{A}^{-1}\,\mathbf{X}\,,$ (54)
where $\mathbb{A}$ is a constant regular square $n\times n$ matrix, and the system (53) becomes
$\dot{\mathbf{Y}}=\mathbb{M}^{\prime}\,\mathbf{Y}\,,\qquad{\rm with}\qquad\mathbb{M}^{\prime}=\mathbb{A}^{-1}\,\mathbb{M}\,\mathbb{A}\,.$ (55)
Since both transformations are linear and finite dimensional, they are continuous, whence it follows that (53) is stable if, and only if, (55) is stable too. Indeed, introducing the norm $\,\|\mathbf{X}\|_{\infty}:=\overline{\rm ext\,}\{|x_{1}|,\ldots|x_{n}|\}\,$, we have that
$\|\mathbf{Y}(t)\|_{\infty}\leq n\,\|\mathbb{A}^{-1}\|_{\infty}\,\|\mathbf{X}(t)\|_{\infty}\qquad{\rm and}\qquad\|\mathbf{X}(t)\|_{\infty}\leq n\,\|\mathbb{A}\|_{\infty}\,\|\mathbf{Y}(t)\|_{\infty}\,.$
The proof follows easily after combining the latter with Definition 1, namely
$\forall\,\varepsilon>0\,,\quad\exists\,\delta\quad\mbox{such that}\quad\|\mathbf{X}(0)\|_{\infty}<\delta\qquad\Rightarrow\qquad\|\mathbf{X}(t)\|_{\infty}<\varepsilon\,,\quad\forall\,t\,.$
Let us choose $\mathbb{A}$ so that $\mathbb{M}^{\prime}$ is the canonical Jordan matrix, which is made of as many diagonal blocks as characteristic roots, that is
[…]
$\mathbb{M}_{\alpha}$ is a square $n_{\alpha}\times n_{\alpha}$ matrix, where $n_{\alpha}$ is the dimension of the corresponding invariant subspace,
$\mathbb{M}_{\alpha}=\left(\begin{array}[]{cccc}\lambda_{\alpha}&&\ldots&\\ \sigma_{1}&\lambda_{\alpha}&&\vdots\\ 0&\sigma_{2}&\lambda_{\alpha}&\\ &\ldots&&\ddots\end{array}\right)\,,\qquad\mbox{where $\sigma_{j}$ is 0 or 1.}$.
Equation (55) then amounts to
$\dot{\mathbf{Y}}_{\alpha}=\mathbb{M}^{\prime}_{\alpha}\,\mathbf{Y}_{\alpha}\,,\quad\forall\,\alpha\,,$
whose general solution is
$\mathbf{Y}_{\alpha}(t)=e^{t\lambda_{\alpha}}\,\mathbf{P}(t)\,,\qquad\mbox{where}\qquad\mathbf{P}(t)=\mathbf{P}_{0}+t\,\mathbf{P}_{1}+t^{2}\,\mathbf{P}_{2}\ldots$ (57)
is a polynomial of degree $n_{\alpha}-1\,$ and its coefficients, $\,\mathbf{P}_{j}\,$, depend linearly on $\,\mathbf{Y}_{\alpha}(0)\,$ up to $n_{\alpha}-1\,$. In case that ${\mathbb{R}}{\rm e}(\lambda_{\alpha})<0\,$, we have that $\,\|\mathbf{Y}_{\alpha}(t)\|\,$ goes to zero when $t\rightarrow\infty\,$.
In case that $\lambda_{\alpha}=i\,\omega_{\alpha}\,$ is imaginary and simple, then (57) is simply
$\mathbf{Y}_{\alpha}(t)=e^{i\,t\omega_{\alpha}}\,\mathbf{Y}_{\alpha}(0)\qquad{\rm and}\qquad\|\mathbf{Y}_{\alpha}(t)\|_{\infty}=\|\mathbf{Y}_{\alpha}(0)\|\,.$
In any case the definition (56) implies that
$\|\mathbf{Y}(t)\|_{\infty}=\overline{\rm ext}\{\|\mathbf{Y}_{\alpha}(t)\|\}$
and therefore, provided that the either characteristic roots have negative real part or are imaginary and simple,
  • either $\,\displaystyle{\lim_{t\rightarrow\infty}\|\mathbf{Y}(t)\|_{\infty}=0}\,$, if no characteristic root is imaginary,
  • or $\,\|\mathbf{Y}(t)\|_{\infty}\leq\|\mathbf{Y}(0)\|_{\infty}\,$, in the oposite case.
The stability of the equilibrium point then follows straightforward.
Appendix B
We shall prove that there are quadratic potentials $\,V=V(q_{a})=\displaystyle{\frac{1}{2}\,\left(a\,q_{1}^{2}+b\,q_{2}^{2}\right)+c\,q_{1}q_{2}}\,$, with $c\neq 0\,$, such that equation (25) admits a solution $\,G(q,p)\,$ that is quadratic and positive. As such, it has a strict minimum at the origin $\,q_{a}=p_{b}=0\,$ and it will serve as a Liapunov function to ensure the stability of the equilibrium point.
The square 4$\times$4 matrices in equation (25) look like:
$\mathbb{H}=\left(\begin{array}[]{c|c}\mathbf{A}&0\\ \hline\cr 0&\mathbf{B}\end{array}\right)\,,\qquad\mathbb{G}=\left(\begin{array}[]{c|c}\mathbf{U}&\mathbf{R}\\ \hline\cr\mathbf{R}^{T}&\mathbf{S}\end{array}\right)\,,\qquad\mathbf{U}=\mathbf{U}^{T}\,,\qquad\mathbf{S}=\mathbf{S}^{T}\,,$ (58)
where
$\mathbf{A}=\left(\begin{array}[]{cc}a+\omega_{1}^{2}&c\\ c&b-\omega_{2}^{2}\end{array}\right)\,,\quad\qquad\mathbf{B}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\,,$ (59)
and the “boldmath” type is used to denote square 2$\times$2 matrices.
Substituting them in the matrix equation (25) and including that $\,\mathbf{A}\,$, $\mathbf{B}\,$, $\mathbf{U}$ and $\mathbf{S}\,$ are symmetric matrices, we obtain:
$\mathbf{A}\,\mathbf{R}^{T}+\mathbf{R}\,\mathbf{A}=0\,,\qquad\mathbf{A}\,\mathbf{S}-\mathbf{U}\,\mathbf{B}=0\,,\qquad\mathbf{B}\,\mathbf{R}+\mathbf{R}^{T}\,\mathbf{B}=0\,.$ (60)
Working these equations we find out that the third one implies that
$\mathbf{R}=\left(\begin{array}[]{cc}0&z\\ z&0\end{array}\right)\,.$
Combining this with the first of equations (60), we obtain that either $z=0$ and $\,\mathbf{A}\,$ is arbitrary or $z\neq 0\,$ and $\,\displaystyle{\mathbf{A}=\left(\begin{array}[]{cc}w&0\\ 0&-w\end{array}\right)}\,,$ which implies a constraint on the input $V\,$. So that, we shall take $\,\mathbf{R}=0\,$ and the integral of motion is
$G=\frac{1}{2}\,U_{ab}\,q_{a}q_{b}+\frac{1}{2}\,S_{ab}\,p_{a}p_{b}\,,$ (61)
As for the second equation (60), it amounts to
$\mathbf{U}=\mathbf{A}\,\mathbf{S}\,\mathbf{B}\,,$ (62)
where we have included that $\,\mathbf{B}^{2}\,$ is unity. Putting $\,w:=a+\omega_{1}^{2}\,,\;\,v:=b-\omega_{2}^{2}\,$, we have
$\mathbf{A}=\left(\begin{array}[]{cc}w&c\\ c&v\end{array}\right)\qquad\quad{\rm and}\qquad\quad\mathbf{S}=\left(\begin{array}[]{cc}r&s\\ s&t\end{array}\right)\,,$
and equation (62) leads to
$\mathbf{U}=\left(\begin{array}[]{cc}rw+cs&-ws-ct\\ rc+vs&-cs-vt\end{array}\right)\,.$ (63)
Since $\mathbf{U}$ is symmetric, we have that
$(w+v)\,s+c\,(t+r)=0\,,$ (64)
which is a constraint on the elements of the matrix $\mathbf{S}\,$.
We have thus determined the three blocks of the matrix $\mathbb{G}$, which must define a positive quadratic form. Since $\mathbf{R}=0\,$, this amounts to require $\mathbf{S}\,$ and $\mathbf{U}\,$ to be both positive, which happens if, and only if,
$r>0\,,\qquad\quad\det(\mathbf{S})>0\,,\qquad\quad w+cs>0\,,\qquad\quad\det(\mathbf{U})>0\,.$ (65)
The second inequality reads
$r\,t>s^{2}\,,$ (66)
which implies that both $r$ and $t$ do not vanish and have the same sign. Hence $\,t+r\neq 0\,$ that, together with (64) and that $\,c\neq 0\,$, implies that $\,v+w\neq 0\,$ and $\,s\neq 0\,$.
Combining (64) and (66), we have $\;\displaystyle{\frac{c^{2}}{(w+v)^{2}}(t+r)^{2}<r\,t\leq\frac{1}{4}\,(t+r)^{2}}\,$, and this implies that
$(w+v)^{2}>4\,c^{2}\,,$ (67)
which is a condition on the coefficients of the potential $\,V\,$.
Combined with (62) and the fact that $\,\det\mathbf{B}=-1\,$, the fourth inequality (65) amounts to $\,\det(\mathbf{A})<0\,$, that is
$w\,v<c^{2}\,.$ (68)
Since we only need a particular solution of (25), we take
$r=t=1\qquad{\rm and}\qquad s=-\frac{2c}{w+v}\,,$ (69)
which does satisfy the constraint (64) and the inequality (66), thanks to the condition (67).
Using this, we have that the third inequality (65) becomes
$w-\frac{2c^{2}}{w+v}>0\,,$ (70)
which can be easily proved to be equivalent to
$w-v>0\,,$ (71)
provided that the inequalities (67) and (68) hold. Indeed, let us assume (70), we have that
If $w+v>0\,$,
then $\,w(w+v)>2c^{2}>2\,wv\,$, where (68) has been included. This implies that: (i) $\;\,w(v+w)>0\,$ and therefore $w$ is positive, and (ii) $\;\,w(w-v)>0\,$, hence $\;w-v>0\,$.
If $w+v<0\,$,
then, using (67), we have that $\;\displaystyle{w(w+v)<2c^{2}<\frac{1}{2}\,(w+v)^{2}}\,$, which leads to $\;\displaystyle{(w-v)(w+v)<0\,}$ and (71) follows immediately.
Conversely, starting from (71), we have that
If $w+v>0\,$,
then $\,w-v>0\,$ implies that $\,w>0\,$ and
$w(w+v)=\frac{1}{2}\,(w+v)^{2}+\frac{1}{2}\,(w^{2}-v^{2})>\frac{1}{2}\,(w+v)^{2}>2\,c^{2}\,,$
whence (70) follows.
If $w+v<0\,$,
then $\,w-v>0\,$ implies that $\,v<0\,$. If $\,w>0\,$, then $\,\displaystyle{w-\frac{2c^{2}}{w+v}=w+\frac{2c^{2}}{|w+v|}>0\,}$,
while if $\,w<0\,$, then $\;\displaystyle{w-\frac{2c^{2}}{w+v}=\frac{2c^{2}-|w|\cdot|w+v|}{|w+v|}>2\,\frac{c^{2}-|w|\cdot|v|}{|w+v|}=2\,\frac{c^{2}-wv}{|w+v|}>0\,}$,
where (68) has been included. $\Box$
The region defined by the inequalities (67), (68) and (71) is the shadowed area in the figure below and it is not empty.
Summarizing, a particular solution to equation (24), such that it is positive except at the origin, exists as long as the coefficients of the potential $\,V(q)\,$ fulfill the inequalities (67), (68) and (71) that, in terms of the coefficients of the potential, can be written as
$(a+b+\omega_{1}^{2}-\omega_{2}^{2})^{2}>4\,c^{2}>4\,(a+\omega_{1}^{2})\,(b-\omega_{2}^{2})\qquad{\rm and}\qquad a+\omega_{1}^{2}>b-\omega_{2}^{2}\,.$ (72)
Substituting (61), we finally obtain
$G=h_{1}+h_{2}+V+\frac{c^{2}q_{1}^{2}-\left[c^{2}+b(a+b+\omega_{1}^{2}-\omega_{2}^{2})\right]\,q_{2}^{2}+2c\left[b-\omega_{2}^{2}\right]q_{1}q_{2}+2cp_{1}p_{2}}{a+b+\omega_{1}^{2}-\omega_{2}^{2}}\,,$ (73)
with $\,w+v=a+b+\omega_{1}^{2}-\omega_{2}^{2}\,$.
Appendix C
To find the solution of equation (29) for the particular case $\,V=V(z)\,$, where $\,z(q_{a})\,$ is some quartic polynomial. We first introduce the new variables
$x=1+2\,\left(q_{1}^{2}+q_{2}^{2}\right)\,,\qquad y=2\,\left(q_{1}^{2}-q_{2}^{2}\right)\,,$
then the partial derivatives transform as
$\partial_{1}=4q_{1}\,\left(\partial_{x}+\partial_{y}\right)\qquad{\rm and}\qquad\partial_{2}=4q_{2}\,\left(\partial_{x}-\partial_{y}\right)\,,$
and equations (29) become
$\partial_{y}W=x\,\partial_{x}V\qquad{\rm and}\qquad\partial_{x}W=\partial_{y}V-y\,\partial_{x}V\,,$ (74)
that is $\quad{\rm d}\,W=\left(\partial_{y}V-y\,\partial_{x}V\right)\,{\rm d}x+x\,\partial_{x}V\,{\rm d}y\,$.
If we look for solutions of the kind $\;V=V(z)\quad{\rm with}\quad z=z(x,y)\,$, the latter equation becomes
${\rm d}\,W=V^{\prime}(z)\,\theta\,,\qquad{\rm with}\qquad\theta=\left(\partial_{y}z-y\,\partial_{x}z\right)\,{\rm d}x+x\,\partial_{x}z\,{\rm d}y\,.$
The integrability condition that follows from $\,{\rm d}^{2}W=0\,$ amounts to $\,\,V^{\prime}(z)\,{\rm d}\theta+V^{\prime\prime}(z)\,{\rm d}z\wedge\theta=0\,$, which implies that it exits a function $\,\gamma(z)\,$ such that
${\rm d}z\wedge\theta=\gamma(z)\,{\rm d}\theta\quad\qquad{\rm and}\qquad\quad V^{\prime}(z)+\gamma(z)\,V^{\prime\prime}(z)=0\,,$ (75)
or
$V^{\prime}(z)=K\cdot e^{-\int\gamma^{-1}\,{\rm d}z}\,,\qquad K={\rm constant}\,.$ (76)
The first equation (75) has many solutions. Particularly if we take for $z$ a quadratic polynomial
$z(x,y)=\frac{a}{2}\,x^{2}+\frac{b}{2}\,y^{2}+c\,xy+e\,x+f\,y+g\,,$
after some work we obtain that the first of equations (62) admits five classes of solutions for $z(x,y)\,$, one purely linear and four genuinely quadratic:
  • I.-
    $\,\displaystyle{z=x+by\quad{\rm and}\quad\gamma=\frac{1}{2}\,\left(z-b^{2}\right)}\,$. Then (75) and (76) lead to
    $V=\frac{C}{x+yb-b^{2}}\qquad{\rm and}\qquad W=\frac{x}{b}\,V\,,\qquad C={\rm constant}$
    or, using the definition of the variables $x$ and $y$,
    $V=\frac{C}{1-b^{2}+2\,q_{1}^{2}(1+b)+2\,q_{2}^{2}(1-b)}\qquad{\rm and}\qquad W=\frac{1}{b}\,\left(1+2q_{1}^{2}+2q_{2}^{2}\right)\,V\,.$ (77)
  • II.-
    $\,\displaystyle{z=\frac{1}{2}\,y^{2}+ay\quad{\rm and}\quad\gamma=2z+a^{2}}$ and, using (75) and (76), we obtain
    $\,V=C(y+a)\quad{\rm and}\quad W=Cx\,$, or
    $V(q)=C\,\left[2\,\left(q_{1}^{2}-q_{2}^{2}\right)+a\right]\quad{\rm and}\quad W(q)=C\,\left[1+2\,\left(q_{1}^{2}+q_{2}^{2}\right)\right]\,.$ (78)
  • III.-
    $\,\displaystyle{z=\frac{1}{2}\,y^{2}+2x\quad{\rm and}\quad\gamma=\frac{2}{3}\,z}\,,$ and we get $\,\displaystyle{V=-\frac{2^{3/2}\,K}{\sqrt{y^{2}+4x}}\quad{\rm and}\quad W=-\frac{y}{2}\,V\,}$, that is
    $V(q)=\frac{C}{\sqrt{1+2\,(q_{1}^{2}+q_{2}^{2})+(q_{1}^{2}-q_{2}^{2})^{2}}}\qquad{\rm and}\qquad W=\left(q_{2}^{2}-q_{1}^{2}\right)V\,.$ (79)
  • IV.-
    $\,\displaystyle{z=\frac{1}{2}\,x^{2}\quad{\rm and}\quad\gamma=\frac{2}{3}\,z}\,,$ whence it follows that $\,\;\displaystyle{V=-\frac{2^{3/2}\,K}{x}\quad{\rm and}\quad W=-y\,V}\,$ or,
    $V(q)=\frac{C}{{1+2\,(q_{1}^{2}+q_{2}^{2})}}\qquad{\rm and}\qquad W(q)=2\,\left(q_{2}^{2}-q_{1}^{2}\right)\,V\,.$ (80)